A metallurgist needs to make 12.4 lb of an alloy containing 50% gold. He is going to melt and combine one metal that is 60% gold with another metal that is 40% gold. How much of each should he use?
Answer by robertb(5830)

0.60x + 0.40(12.4 - x) = 0.50*12.4,
0.60x + 4.96 - 0.40x = 6.2,
0.20x = 1.24,
x = 6.2. He has to mix 6.2 lb of the 60% gold alloy, and the same amount of the 40% gold alloy.